A father pushes his young daughter on a sled up a 15o inclined-plane which is moving at a


constant velocity. If the push of the father is directed horizontally with a magnitude of


92.0 N, (a) how much work is done by the applied force of the father on the sled with the girl


up to a height of 3.25 m from the base of the inclined plane? (b) What is the work done by


the weight on the sled with the girl if its total mass is 35.0 kg?



1
Expert's answer
2022-02-16T08:35:08-0500

Given:

"\theta=15^\circ"

"F=92.0\rm \: N"

"h=3.25\:\rm m"

"m=35.0\:\rm kg"


(a) the work done by the applied force

"W_1=Fd=Fh/\sin\theta\\
=\rm 92.0*3.25/\sin15^\circ=1155\: J"

(b) the work done by the gravity force

"W_2=mgh\\
=35.0*9.8*3.25=1115\:\rm J"